OpenStudy (anonymous):

Please help, it's an easy tangent line equation question: Find the equation to the tangent line to the curve f(x)= (x+3)^(1/5), at x=3

OpenStudy (anonymous):

\[f(x)=\sqrt[5]{x+3}\]

OpenStudy (zarkon):

find the derivative

OpenStudy (anonymous):

I tried, but it gave me weird things... the square root confused me

OpenStudy (zarkon):

\[\frac{d}{dx}(f(x))^{1/n}=\frac{1}{n}(f(x))^{\frac{1}{n}-1}f'(x)\]

OpenStudy (zarkon):

also it is not a square root...it is a 5th root

OpenStudy (anonymous):

Thank you!

OpenStudy (zarkon):

np